OpenStudy (anonymous):

Use a graph and a table to solve the system {3y+6x=3 x-y=7 and i'm suppose to have the two points where they intercept but i don't know how to do it.

OpenStudy (mertsj):

Divide the first equation by 3

OpenStudy (mertsj):

You will get 2x+y=1 Then add the two equations. That will eliminate y and you can solve for x. can you do that?

OpenStudy (anonymous):

im still a little confused..

OpenStudy (anonymous):

no. it says graph and table!

OpenStudy (anonymous):

Graph/Plot 3y + 6x = 3 Take 3y + 6x = 3 and turn it into slope intercept form 3y = -6x +3 3y/3 = -6x/3 +3/3 y = -2x + 1 in this equation we have our first point y = -2x + 1, which is the y intercept 1. plot (0,1) Now we have our first point (0,1) lets use our slope to find the next point. Our slope is -2x, so from (0,1) move down 2, since we are working with negatives, form the point (0,1), which will give us (1,-1).Now plot the two points (0,1) and (1,-1) and draw a line. do the same for the other equation and where the two lines cross is your answer. This is if you have to graph your work and show it

OpenStudy (anonymous):

graph the 2 equations. the first one is 3y+6x=3 --> y+2x =1 --> y=1-2x the second one is x-y=7 --> x = 7+y --> y = x-7

OpenStudy (anonymous):

good job doodletech

OpenStudy (mertsj):